# Optimized neural forms for solving ordinary differential equations

### v1

@@ 8 Conclusions
Note that since neural methods rely on multidimensional optimization, careful selection of the employed solvers is important. Commonly utilized methods crafted for big data, such as stochastic gradient descent, Adam, and their variants, may not perform well. Instead, pattern search and quasi-Newton methods were found to be very effective. In addition, we have noticed that training datasets based on Chebyshev grids, were superior to those based on equidistant grids, for the majority of the examined test problems. This advantage stems from their ability to mitigate the Runge phenomenon which is related to overtraining.
Overall, the proposed approach is easy to implement and offers a differentiable closed-form solution that (i) has superior interpolation capabilities, and (ii) can be directly employed in subsequent related calculations. Deep or shallow neural networks, with any valid activation function (that meets the universal approximation requirements) may be used.
Fertile ground for future work is offered for extending the proposed ideas. An interesting path would be the use of the recently introduced functionally weighted neural network, which is a special type of feedforward neural network with advanced extrapolation capabilities [13, 18]. Moreover, the proposed ideas may be extended, in a rather straightforward fashion, to treat challenging stiff ODEs, and PDEs in rectangular domains. Further elaboration and research is required for solving PDEs defined inside non-rectangular domains.
Appendix A Results for the Runge-Kutta method
| Test | Condition | Number of points | ||||||||
| Problem | Type | Training dataset | Test dataset | |||||||
| 1 | 40 | 80 | 160 | 1000 | ||||||
| MSD | MXD | MSD | MXD | MSD | MXD | MSD | MXD | |||
| D | 0.61e-03 | 0.20e-01 | 0.52e-06 | 0.22e-04 | 0.90e-09 | 0.37e-07 | 0.57e-15 | 0.26e-13 | ||
| 2 | 90 | 180 | 270 | 1000 | ||||||
| MSD | MXD | MSD | MXD | MSD | MXD | MSD | MXD | |||
| D-D | 0.46e-09 | 0.81e-08 | 0.24e-11 | 0.43e-10 | 0.10e-12 | 0.18e-11 | 0.32e-17 | 0.59e-16 | ||
| D-N | 0.42e-09 | 0.34e-07 | 0.14e-11 | 0.17e-09 | 0.49e-13 | 0.74e-11 | 0.11e-17 | 0.23e-15 | ||
| N-N | 0.42e-09 | 0.34e-07 | 0.14e-11 | 0.17e-09 | 0.49e-13 | 0.74e-11 | 0.11e-17 | 0.23e-15 | ||
| C | 0.12e-08 | 0.86e-07 | 0.39e-11 | 0.43e-09 | 0.14e-12 | 0.18e-10 | 0.36e-17 | 0.56e-15 | ||
| R | 0.33e-09 | 0.25e-07 | 0.11e-11 | 0.13e-09 | 0.39e-13 | 0.54e-11 | 0.93e-18 | 0.17e-15 | ||
| 3 | 180 | 360 | 720 | 1000 | ||||||
| MSD | MXD | MSD | MXD | MSD | MXD | MSD | MXD | |||
| C | 0.11e-07 | 0.68e-07 | 0.65e-09 | 0.42e-08 | 0.41e-10 | 0.26e-09 | 0.11e-10 | 0.70e-10 | ||
| 4 | 70 | 130 | 250 | 1000 | ||||||
| MSD | MXD | MSD | MXD | MSD | MXD | MSD | MXD | |||
| D | 0.15e-09 | 0.25e-08 | 0.87e-12 | 0.16e-10 | 0.42e-14 | 0.78e-13 | 0.59e-19 | 0.11e-17 | ||
| D | 0.10e-08 | 0.20e-07 | 0.61e-11 | 0.13e-09 | 0.29e-13 | 0.63e-12 | 0.41e-18 | 0.90e-17 | ||
| 5 | 70 | 130 | 250 | 1000 | ||||||
| MSD | MXD | MSD | MXD | MSD | MXD | MSD | MXD | |||
| D | 0.70e-04 | 0.37e-02 | 0.18e-06 | 0.85e-05 | 0.51e-09 | 0.26e-07 | 0.47e-14 | 0.24e-12 | ||
| D | 0.16e-03 | 0.82e-02 | 0.39e-06 | 0.19e-04 | 0.11e-08 | 0.58e-07 | 0.11e-13 | 0.54e-12 | ||
| 6 | 180 | 270 | 360 | 1000 | ||||||
| MSD | MXD | MSD | MXD | MSD | MXD | MSD | MXD | |||
| D-D | 0.24E-11 | 0.43E-10 | 0.10E-12 | 0.18E-11 | 0.11E-13 | 0.19E-12 | 0.32E-17 | 0.59E-16 | ||
| D-N | 0.14E-11 | 0.17E-09 | 0.49E-13 | 0.74E-11 | 0.46E-14 | 0.77E-12 | 0.11e-17 | 0.23e-15 | ||
| N-N | 0.14E-11 | 0.17E-09 | 0.49E-13 | 0.74E-11 | 0.46E-14 | 0.77E-12 | 0.11e-17 | 0.23e-15 | ||
| C | 0.39E-11 | 0.43E-09 | 0.14E-12 | 0.18E-10 | 0.14E-13 | 0.19E-11 | 0.36e-17 | 0.56e-15 | ||
| 7 | 80 | 160 | 320 | 1000 | ||||||
| MSD | MXD | MSD | MXD | MSD | MXD | MSD | MXD | |||
| D | 0.65e-09 | 0.31e-08 | 0.24e-11 | 0.12e-10 | 0.28e-13 | 0.16e-12 | 0.19e-13 | 0.87e-13 | ||
| Abbreviations: D=Dirichlet, N=Neumann, D-N=Mixed Dirichlet-Neumann, R=Robin, C=Cauchy | ||||||||||
@@ Appendix B Results for the baseline neural method (standalone neural network)
| Measure: Deviation Grid type: Equidistant | ||||||||||||
| Network Parameters | ||||||||||||
| 72 | 144 | 288 | ||||||||||
| Training | Test | Training | Test | Training | Test | |||||||
| MSD | MXD | MSD | MXD | MSD | MXD | MSD | MXD | MSD | MXD | MSD | MXD | |
| 80 | 0.19e-13 | 0.86e-13 | 0.19e-13 | 0.87e-13 | ||||||||
| 160 | 0.19e-13 | 0.87e-13 | 0.19e-13 | 0.87e-13 | 0.19e-13 | 0.87e-13 | 0.19e-13 | 0.87e-13 | ||||
| 320 | 0.19e-13 | 0.87e-13 | 0.19e-13 | 0.87e-13 | 0.19e-13 | 0.87e-13 | 0.19e-13 | 0.87e-13 | 0.19e-13 | 0.87e-13 | 0.19e-13 | 0.87e-13 |
| Measure: Deviation Grid type: Chebyshev | ||||||||||||
| Network Parameters | ||||||||||||
| 72 | 144 | 288 | ||||||||||
| Training | Test | Training | Test | Training | Test | |||||||
| MSD | MXD | MSD | MXD | MSD | MXD | MSD | MXD | MSD | MXD | MSD | MXD | |
| 80 | 0.23e-13 | 0.87e-13 | 0.19e-13 | 0.87e-13 | ||||||||
| 160 | 0.23e-13 | 0.87e-13 | 0.19e-13 | 0.87e-13 | 0.23e-13 | 0.87e-13 | 0.19e-13 | 0.87e-13 | ||||
| 320 | 0.23e-13 | 0.87e-13 | 0.19e-13 | 0.87e-13 | 0.23e-13 | 0.87e-13 | 0.19e-13 | 0.87e-13 | 0.23e-13 | 0.87e-13 | 0.19e-13 | 0.87e-13 |
| Measure: Error Grid type: Equidistant | ||||||||||||
| Network Parameters | ||||||||||||
| 72 | 144 | 288 | ||||||||||
| Training | Test | Training | Test | Training | Test | |||||||
| MSE | MXE | MSE | MXE | MSE | MXE | MSE | MXE | MSE | MXE | MSE | MXE | |
| 80 | 0.41e-20 | 0.18e-19 | 0.14e-19 | 0.22e-17 | ||||||||
| 160 | 0.24e-20 | 0.27e-19 | 0.25e-20 | 0.27e-19 | 0.70e-21 | 0.39e-20 | 0.80e-21 | 0.27e-19 | ||||
| 320 | 0.75e-20 | 0.33e-19 | 0.75e-20 | 0.33e-19 | 0.17e-20 | 0.60e-19 | 0.16e-20 | 0.68e-19 | 0.15e-21 | 0.22e-20 | 0.15e-21 | 0.24e-20 |
| Measure: Error Grid type: Chebyshev | ||||||||||||
| Network Parameters | ||||||||||||
| 72 | 144 | 288 | ||||||||||
| Training | Test | Training | Test | Training | Test | |||||||
| MSE | MXE | MSE | MXE | MSE | MXE | MSE | MXE | MSE | MXE | MSE | MXE | |
| 80 | 0.54e-20 | 0.23e-19 | 0.57e-20 | 0.24e-19 | ||||||||
| 160 | 0.13e-19 | 0.60e-19 | 0.12e-19 | 0.61e-19 | 0.38e-21 | 0.51e-20 | 0.38e-21 | 0.59e-20 | ||||
| 320 | 0.69e-20 | 0.38e-19 | 0.76e-20 | 0.40e-19 | 0.99e-21 | 0.63e-20 | 0.83e-21 | 0.64e-20 | 0.17e-21 | 0.18e-20 | 0.16e-21 | 0.18e-20 |
Appendix C Results for the proposed neural forms
@@
| Measure: Deviation Grid type: Equidistant | ||||||||||||
| Network Parameters | ||||||||||||
| 72 | 144 | 288 | ||||||||||
| Training | Test | Training | Test | Training | Test | |||||||
| MSD | MXD | MSD | MXD | MSD | MXD | MSD | MXD | MSD | MXD | MSD | MXD | |
| 80 | 0.19e-13 | 0.86e-13 | 0.19e-13 | 0.87e-13 | ||||||||
| 160 | 0.19e-13 | 0.87e-13 | 0.19e-13 | 0.87e-13 | 0.19e-13 | 0.87e-13 | 0.19e-13 | 0.87e-13 | ||||
| 320 | 0.19e-13 | 0.87e-13 | 0.19e-13 | 0.87e-13 | 0.19e-13 | 0.87e-13 | 0.19e-13 | 0.87e-13 | 0.19e-13 | 0.87e-13 | 0.19e-13 | 0.87e-13 |
| Measure: Deviation Grid type: Chebyshev | ||||||||||||
| Network Parameters | ||||||||||||
| 72 | 144 | 288 | ||||||||||
| Training | Test | Training | Test | Training | Test | |||||||
| MSD | MXD | MSD | MXD | MSD | MXD | MSD | MXD | MSD | MXD | MSD | MXD | |
| 80 | 0.23e-13 | 0.87e-13 | 0.19e-13 | 0.87e-13 | ||||||||
| 160 | 0.23e-13 | 0.87e-13 | 0.19e-13 | 0.87e-13 | 0.23e-13 | 0.87e-13 | 0.19e-13 | 0.87e-13 | ||||
| 320 | 0.23e-13 | 0.87e-13 | 0.19e-13 | 0.87e-13 | 0.23e-13 | 0.87e-13 | 0.19e-13 | 0.87e-13 | 0.23e-13 | 0.87e-13 | 0.19e-13 | 0.87e-13 |
| Measure: Error Grid type: Equidistant | ||||||||||||
| Network Parameters | ||||||||||||
| 72 | 144 | 288 | ||||||||||
| Training | Test | Training | Test | Training | Test | |||||||
| MSE | MXE | MSE | MXE | MSE | MXE | MSE | MXE | MSE | MXE | MSE | MXE | |
| 80 | 0.23e-19 | 0.13e-18 | 0.77e-19 | 0.82e-17 | ||||||||
| 160 | 0.13e-19 | 0.51e-19 | 0.13e-19 | 0.52e-19 | 0.19e-20 | 0.20e-19 | 0.70e-20 | 0.17e-17 | ||||
| 320 | 0.91e-20 | 0.17e-18 | 0.87e-20 | 0.17e-18 | 0.20e-20 | 0.23e-19 | 0.19e-20 | 0.25e-19 | 0.12e-20 | 0.29e-19 | 0.12e-20 | 0.30e-19 |
| Measure: Error Grid type: Chebyshev | ||||||||||||
| Network Parameters | ||||||||||||
| 72 | 144 | 288 | ||||||||||
| Training | Test | Training | Test | Training | Test | |||||||
| MSE | MXE | MSE | MXE | MSE | MXE | MSE | MXE | MSE | MXE | MSE | MXE | |
| 80 | 0.15e-19 | 0.63e-19 | 0.18e-19 | 0.69e-19 | ||||||||
| 160 | 0.56e-19 | 0.31e-18 | 0.69e-19 | 0.34e-18 | 0.16e-20 | 0.60e-20 | 0.17e-20 | 0.71e-20 | ||||
| 320 | 0.58e-20 | 0.45e-19 | 0.55e-20 | 0.47e-19 | 0.20e-20 | 0.10e-19 | 0.20e-20 | 0.11e-19 | 0.27e-20 | 0.17e-19 | 0.21e-20 | 0.18e-19 |

### v2

@@ 8 Conclusions
Moreover, our analysis indicates that the careful selection of optimizer for training the employed neural networks may offer advantages. Commonly utilized methods crafted for big data applications, such as stochastic gradient descent, Adam, and variants, may not perform well. Instead, pattern search and quasi-Newton methods were found to be very effective. In addition, we have noticed that Chebyshev grids performed better than equidistant grids for the majority of the considered test problems. This advantage stems from their capability to mitigate the Runge phenomenon which is related to overtraining of the neural networks.
Overall, the proposed approach is easy to implement and offers a differentiable closed-form solution that can be directly employed in subsequent calculations. Deep or shallow neural networks with any activation function meeting the universal approximation requirements may be used. The proposed ideas have the potential to provide a solid foundation for future initiatives. It might be of interest to consider the recently introduced functionally weighted neural network [13, 18], which is a special type of feedforward neural network with advanced extrapolation capabilities. Furthermore, the proposed ideas could be extended in a straightforward manner to treat PDEs in rectangular domains. The challenging problem of PDEs defined inside non-rectangular domains certainly merits attention and requires further investigation.
Appendix A Results for the ode113 solver
| Test | Condition | Number of | Spline Polynomial | |||||
| Problem | Type | Training Points | Degree | Training points | Test dataset | |||
| 1 | MSD | MXD | MSD | MXD | ||||
| D | 300 | 3 | 0.58e-28 | 0.18e-26 | 0.12e-16 | 0.76e-14 | ||
| 5 | 0.12e-23 | 0.99e-21 | ||||||
| 7 | 0.15e-13 | 0.15e-10 | ||||||
| 3 | MSD | MXD | MSD | MXD | ||||
| C | 215 | 3 | 0.31e-29 | 0.28e-28 | 0.14e-15 | 0.18e-14 | ||
| 5 | 0.12e-16 | 0.73e-14 | ||||||
| 7 | 0.70e+18 | 0.33e+21 | ||||||
| 4 | MSD | MXD | MSD | MXD | ||||
| D | 109 | 3 | 0.20e-24 | 0.51e-23 | 0.56e-16 | 0.27e-15 | ||
| 0.14e-23 | 0.41e-22 | 0.19e-23 | 0.41e-22 | |||||
| 5 | 0.91e-24 | 0.63e-22 | ||||||
| 0.20e-23 | 0.75e-22 | |||||||
| 7 | 0.10e-08 | 0.54e-06 | ||||||
| 0.80e-06 | 0.42e-03 | |||||||
| 5 | MSD | MXD | MSD | MXD | ||||
| D | 321 | 3 | 0.52e-29 | 0.86e-28 | 0.97e-17 | 0.14e-14 | ||
| 0.16e-28 | 0.31e-27 | 0.22e-16 | 0.31e-14 | |||||
| 5 | 0.72e-24 | 0.37e-21 | ||||||
| 0.16e-23 | 0.84e-21 | |||||||
| 7 | 0.75e-09 | 0.59e-06 | ||||||
| 0.16e-11 | 0.13e-08 | |||||||
| 6 | MSD | MXD | MSD | MXD | ||||
| D-D | 203 | 3 | 0.13e-26 | 0.56e-26 | 0.11e-16 | 0.34e-14 | ||
| 5 | 0.63e-22 | 0.52e-19 | ||||||
| 7 | 0.35e-16 | 0.27e-13 | ||||||
| D-N | 203 | 3 | 0.11e-26 | 0.46e-26 | 0.11e-16 | 0.34e-14 | ||
| 5 | 0.63e-22 | 0.52e-19 | ||||||
| 7 | 0.35e-16 | 0.26e-13 | ||||||
| N-N | 203 | 3 | 0.42e-27 | 0.14e-26 | 0.11e-16 | 0.34e-14 | ||
| 5 | 0.63e-22 | 0.52e-19 | ||||||
| 7 | 0.52e-16 | 0.40e-13 | ||||||
| C | 203 | 3 | 0.27e-24 | 0.42e-23 | 0.11e-16 | 0.34e-14 | ||
| 5 | 0.64e-22 | 0.53e-19 | ||||||
| 7 | 0.41e-16 | 0.31e-13 | ||||||
| Abbreviations: D=Dirichlet, N=Neumann, D-N=Mixed Dirichlet-Neumann, C=Cauchy | ||||||||
@@ Appendix B Results for the baseline neural method (standalone neural network)
Appendix C Results for the proposed augmented neural forms